Q.1: Apply Gauss's Divergence theorem to evaluate double integration S (lx^ 2 +my^ 2 +nz^ 2 )dS taken over the sphere (x-a)^ 2 +(y-b)^ 2 +(z-c)^ 2 = rho^ 2 ;l,m,n being the direction cosines of the external normal to the sphere.

1
Expert's answer
2021-11-01T12:10:13-0400

By Gauss divergence theorem\\"\iint_s\bar{F.\bar{n}}\space ds=\iiint_v\space div\space dv"

where "\bar{F}.\bar{n}=lx^2+my^2+nz^2"

let "\bar{n}=li+mj+nk\space then \space \bar{F}=xi^2+y@j+n62\bar{k}"

so that div "\bar{F}=\frac{\delta}{\delta x}(x^2)+\frac{\delta}{\delta y}(y^2)+\frac{\delta}{\delta z}(z^2)"

"=2x+2y+2z\\=2(x+y+z)"

"\iiint_v div \bar{F} div=\iiint_v 2(x+y+z)dxdydz....(1)"

now consider "\iiint_v 2x\space dx\space dy\space dz"

use spherical coordinate

"x-a=sin \theta\space cos\theta\\

y-b=sin \theta\space sin\theta\\z-c=cos\theta"

"dx\space dy\space dz=sin \theta\space cos \theta"

and "0≤\theta≤\pi,0=\theta≤2\pi,0≤e"

"\iiint_v 2x\space dx\space dy\space dz=\int^{\pi }_0 \int ^{2\pi}_0 \int ^e_0 2(a+r sin \theta cos \theta)r^2 sin\theta dvd \theta d \theta"


"\int^{\pi }_0 \int ^{2\pi}_0 \int ^e_0 2ar^2+ sin \theta dvd \theta d \theta+
\int^{\pi }_0 \int ^{2\pi}_0 \int ^e_0 r^3+ sin ^2\theta dvd \theta d \theta"

"=-cos\int ^{\pi}_0 (2ae^3)\int ^{2\pi}_0 d\theta+{\int\pi}_0(\frac{1-cos \theta}{2}) d\theta(sin\theta)\int^{2\pi}_0(\frac{e^4}{4})"

"=2\frac{2ae^3}{3}(2\pi-0)\\=\frac{8}{3}a\pi e^3"

similarly"\iiint 2y\space dx\space dy\space dz=\frac{8}{3}b \pi e^3"

and "\iiint 2z\space dx\space dy\space dz=\frac{8}{3}c \pi e^3"

by adding these three integrals we get

"\iint_s\bar{F.\bar{n}}\space ds=\iiint_v\space div\space \bar{F} dv=8\pi e^3(a+b+c)"